PLEASE HELP!
1. Multiply (c+ 1/2) (c+1/5)

2.Factor c^2+10c+25 = My answer is (c-5)(c+5)?
THANK YOU, THANK YOU, THANK YOU

Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

1. Multiply (c+ 1/2) (c+1/5)
c(c+1/5)+1/2(c+1/5)
c^2+c/5 +c/2 + 1/10
c^2+7c/10 +1/10

2.Factor c^2+10c+25 = My answer is (c-5)(c+5)?
This is an identity.
a^2+2ab+b^2=(a+b)^2
c^2+2*1*5c+5^2
(c+5)^2

Answer by stanbon(75887)

1. Multiply (c+ 1/2) (c+1/5)
Use FOIL:
F: c^2
O:c/5
I:c/2
L:1/10
---
Combine to get:
c^2 + (1/2 + 1/5)c + 1/10
---
c^2 + (7/10)c + 1/10
===============================
2.Factor c^2+10c+25
(c+5)(c+5)
= (c+5)^2
===================
Cheers,
Stan H.